### v3

# A BENCHMARK AND EVALUATION FOR REAL-WORLD OUT-OF-DISTRIBUTION DETECTION USING VISION-LANGUAGE MODELS

Abstract
Out-of-distribution (OOD) detection is a task that detects OOD samples during inference to ensure the safety of deployed models. However, conventional benchmarks have reached performance saturation, making it difficult to compare recent OOD detection methods. To address this challenge, we introduce three novel OOD detection benchmarks that enable a deeper understanding of method characteristics and reflect real-world conditions. First, we present ImageNet-X, designed to evaluate performance under challenging semantic shifts. Second, we propose ImageNet-FS-X for full-spectrum OOD detection, assessing robustness to covariate shifts (feature distribution shifts). Finally, we propose Wilds-FS-X, which extends these evaluations to real-world datasets, offering a more comprehensive testbed. Our experiments reveal that recent CLIP-based OOD detection methods struggle to varying degrees across the three proposed benchmarks, and none of them consistently outperforms the others. We hope the community goes beyond specific benchmarks and includes more challenging conditions reflecting real-world scenarios. The code is https://github.com/hoshi23/OOD-X-Benchmarks.
Index Terms—  Out-of-distribution Detection, Vision Language Model, Benchmark
1 Introduction
Machine learning models, including vision-language models (VLMs), have recently improved image recognition but may fail with unknown examples, risking real-world applications [1, 2]. Out-of-distribution (OOD) detection, which identifies samples that deviate from the semantics distribution of the in-distribution (ID) data, is crucial for ensuring model reliability [3]. Recently, VLM-based methods like CLIP [4], have attracted attention, inspiring many approaches in this field [5, 6].
Various methods have been proposed, and high OOD detection performance has been achieved. However, conventional benchmarks have reached performance saturation, making it difficult to distinguish between different approaches. Commonly used evaluation benchmarks include using ImageNet-1k [7] as ID data and other datasets like iNaturalist [8] or Texture [9] as OOD data. These benchmarks have served as widely adopted baselines for comparing the OOD detection performance [5, 6, 10, 11, 12, 13, 14, 15]. However, the large distribution shift between ID and OOD data makes detection easier, making it difficult to analyze and compare the characteristics of different methods.
Fig. 1: Data examples of ID and OOD in ImageNet-X, a benchmark for challenging semantic shifts.
(a) ImageNet-FS-X.
(b) Wilds-FS-X (iWildCam).
(c) Wilds-FS-X (FMoW).
Fig. 2: Data structure and examples in the proposed benchmarks incorporating covariate shifts (feature distribution shifts).
In this paper, we propose three novel benchmarks, supporting a precise understanding of method characteristics and simulations reflecting real-world conditions (Fig. 1, Fig. 2). For challenging and practical evaluations, three critical perspectives are crucial: (1) the semantic similarity between ID and OOD data, (2) two distribution shifts - semantic shift (class shift) and covariate shift (feature distribution shift), and (3) the dataset’s alignment with real-world scenarios. Here, we define label shifts as ‘semantic shifts,’ and changes in image style or quality as ‘covariate shifts’ following [16]. Research addressing both as mentioned in (2) is termed full-spectrum OOD detection [16]. These perspectives reflect real-world scenarios and enable us to conduct more robust evaluations across different OOD detection methods.
Based on the above perspectives, we propose: (B1) ImageNet-X, a benchmark with a small semantic shift, It is created by splitting ID and OOD data from ImageNet-1k [7], considering the hierarchical structure of ImageNet to ensure that the semantic similarity between ID and OOD data is small. (B2) ImageNet-FS-X, a benchmark that incorporates covariate shift into (B1), allowing for a systematic analysis of these two distribution shifts. (B3) Wilds-FS-X, a benchmark that retains the characteristics of (B1) and (B2) while bringing the data closer to real-world scenarios. This benchmark is based on the Wilds [17] and reflects distribution shifts that naturally occur in real-world application scenarios. Evaluating methods using these three progressive datasets enables rigorous analysis under complex, real-world conditions.
In our experiments, we evaluate representative CLIP-based OOD detection methods. Evaluation using the three proposed benchmarks revealed the following insights: (B1) Comparing with the results of the conventional benchmarks, a change in the method performance ranking is observed, and no single winner emerges as the best across all distribution shifts. (B2) CLIP-based OOD detection performance decreases due to covariate shifts, but the method performance ranking remains consistent with constant semantic shift. (B3) Both classification and OOD detection performance are low, making it a highly challenging benchmark, and even though few-shot improved classification accuracy, it does not necessarily improve OOD detection performance.
Major contributions of the paper are:
  • We proposed three benchmarks for challenging and practical evaluations, enabling a precise understanding of method characteristics and simulations of real-world conditions.
  • Our experiments revealed performance variations across benchmarks and significant drops when incorporating covariate shifts and real-world scenarios, highlighting the need for improvement.
2 Benchmark Construction
This section outlines the background and details the construction methods of each of the three proposed benchmarks.
2.1 ImageNet-X
Background. For OOD detection, conventional benchmarks have primarily evaluated the detection of large distribution shifts, including both semantic and covariate shifts (common OOD detection) with ImageNet [18, 19]. In contrast, hard OOD detection has recently been proposed as a problem setting focusing on detecting small semantic shifts between labels. For example, some benchmarks use ImageNet-1k as ID data and use tailored datasets [20, 21] as OOD data [22]. However, using different datasets for ID and OOD may unintentionally incorporate covariate shifts. OOD detection performance is more sensitive to covariate shifts than semantic shifts [23]. Therefore, utilizing both ID and OOD data from a common dataset is considered an effective approach to minimize the impact of covariate shifts, enabling a more rigorous evaluation of semantic shifts.
Additionally, ImageNet-10 and ImageNet-20 [5], as well as ImageNet-Protocol [24], are benchmarks that use datasets split from ImageNet. However, the former has a limited number of classes, and the latter suffers from a considerable semantic gap between ID and OOD, resulting in performance saturation for both. Therefore, a dataset with more classes and guaranteed similarity in semantic shifts is necessary for more accurate performance comparisons.
Proposed Benchmark. We propose ImageNet-X for large-scale and more rigorous evaluation of challenging semantic shifts. In ImageNet-X, we propose a splitting method for ID and OOD labels based on their semantic structure, dividing closely related labels within that structure. For example, we treat “dalmatian” as ID and “Great Pyrenees” as OOD (Fig. 1). We use the WordNet hierarchy [25] to represent the labels’ semantic structure, with each ImageNet label corresponding to a parent class. By using the parent class data, which is the immediate upper-level hierarchy provided on the official ImageNet site [7], the 1000 classes can be grouped into 558 parent classes. This allows for the separation of labels that are particularly closely related.
The splitting process is conducted as follows: First, the labels within each parent class are split into two halves. Next, the ID and OOD labels are balanced by integrating the divided labels, resulting in an equal number of labels for ID and OOD. The number of split classes and samples are shown in Table 1.
2.2 ImageNet-FS-X
Background. An existing problem setting for OOD detection with covariate shifts is full-spectrum OOD (FS-OOD) [16, 22]. In FS-OOD, the ID data is constructed by adding data with the same semantics but different covariate distributions (covariate-shifted ID) to the training data (training ID). When using ImageNet-1k as the training ID, variant datasets (e.g., ImageNet-V2 [26], ImageNet-R [27], and ImageNet-C [28]) are used as covariate-shifted IDs (ImageNet-FS). Covariate-shifted ID is added only during evaluation to assess the performance change of ID data due to covariate shift. However, differences in covariate distributions between ID and OOD data are not considered, making it difficult to systematically evaluate semantic and covariate shifts separately. Moreover,  [29] proposed a benchmark that allows flexible evaluation of semantic shift and covariate shift in OOD data and analyzed their respective impacts. However, this does not account for covariate shifts in ID data and does not cover the FS-OOD problem setting.
Proposed Benchmark. We propose ImageNet-FS-X, which incorporates covariate shift into ImageNet-X (Fig. 2). This involves adding data with the same labels but different covariate distributions to ImageNet-1k and applying the label splitting method from ImageNet-X to construct ID and OOD data.
The data in ImageNet-FS, specifically the training and variant datasets, are separated into ID and OOD data based on their labels. The ID data from the training dataset is used as the training ID, and the ID data from the variant datasets is used as the covariate-shifted ID. On the other hand, the other data is collectively used as OOD data. The number of split classes and samples for each derived dataset are shown in Table 1. This approach aligns the covariate distribution of OOD data with ID data, enabling a more rigorous evaluation of semantic shift and the impact of covariate shifts on both ID and OOD data. The ImageNet-FS-X benchmark allows for a comprehensive evaluation that considers both sensitivity to semantic shifts and robustness to covariate shifts.
Table 1: Data composition of each benchmark.
The number of classes and samples (used in this experiment) assigned to ID and OOD for each source dataset are shown.
#Classes #Samples
Source ID OOD ID OOD
ImageNet-X ImageNet-1k 500 500 25,000 25,000
ImageNet- FS-X ImageNet-V2 500 500 5,000 5,000
ImageNet-R 495 492 14,755 15,245
ImageNet-C 315 313 5,380 4,620
Wilds-FS-X iWildCam Test-ID 91 91 3,801 4,353
iWildCam Test 91 91 14,526 28,265
FMoW Test-ID 31 31 5,109 6,218
FMoW Test 31 31 9,601 12,507
2.3 Wilds-FS-X
Background. An existing benchmark using real-world scenario datasets is WildCapture [30], which evaluates using a dataset of wildlife trap images. However, it does not explicitly separate the covariate distribution differences between the training and evaluation data, and thus, robustness to covariate shifts has not been evaluated.
Proposed Benchmark. Finally, we propose Wilds-FS-X, a benchmark that brings the problem setting of ImageNet-FS-X closer to real-world scenarios by utilizing Wilds [17]. Wilds is a benchmark that reflects covariate shifts naturally occurring in real-world scenarios. This benchmark evaluates task performance by using test data (Test) that shares the same labels as the training data (Training) but differs in domains such as shooting environments. Wilds aims to evaluate the impact of covariate shifts on recognition performance. To facilitate performance comparison, a test dataset with identical covariate distributions to the training data (Test-ID) is also included.
The model is trained using the ID data from Training. For evaluation, the ID data from Test-ID is defined as training ID, the ID data from Test is defined as covariate-shifted ID, and the remaining data from Test-ID and Test are defined as OOD.
Wilds consists of datasets spanning 10 different domains. For this benchmark, we utilized: iWildCam and FMoW.
iWildCam is a dataset of wildlife photos taken by camera traps. Semantic shift corresponds to different animal species, and covariate shift corresponds to changes in camera locations of camera photos (Fig. 2). The labels comprise 182 classes, including 181 animal species, and an additional class labeled “empty” for images containing no animals. Among these, 91 animal labels were randomly selected as ID data, while the remaining 90 animal labels, along with the “empty” class, were used as OOD data. FMoW is a RGB satellite image dataset. Semantic shift corresponds to the building or land use categories, and covariate shift corresponds to changes in the year when the images were captured (Fig. 2). The dataset includes 62 facility types. From these, 31 classes were randomly chosen as ID data, while the remaining 31 classes were designated as OOD data. The specific sample sizes for the splits in this experiment are shown in Table 1. Using these datasets, we can evaluate the detection of fine-grained semantic shifts and robustness to covariate shifts required in real-world scenarios.
3 Experiments
3.1 Baseline Methods
We examine representative zero-shot and few-shot CLIP-based OOD detection methods. Zero-shot and few-shot settings are the most common settings for CLIP-based OOD detection.
Zero-shot OOD detection. Zero-shot OOD detection is a method that does not require additional training with ID data [31]. Zero-shot methods can be broadly categorized into two types. The first type does not use an OOD prompt and relies solely on the outputs of each ID class label from CLIP to calculate OOD scores. We utilize MCM [5] and GL-MCM [10] in our experiments. The second type utilizes OOD prompts, which involve additional labels or prompts specifically designed for OOD detection. We utilize NegLabel [11], EOE [12], and CLIPN [13] in our experiments.
Few-shot OOD detection. Few-shot OOD detection refers to the use of ID images during both the training and inference phases [31] We utilize several methods in which the context of prompts is learned as learnable parameters from few-shot ID data. Few-shot methods are also divided into approaches that use OOD prompts and those that do not. For methods without OOD prompts, we use CoOp [32] and LoCoOp [6]. The hyperparameters are set according to previous studies, with training performed using 16 shots in all cases. The OOD score calculation for these methods can apply the same approach as zero-shot methods, such as MCM and GL-MCM. In this experiment, following the respective papers, CoOp is evaluated using MCM, and LoCoOp is evaluated using GL-MCM. For few-shot methods that use OOD prompts, we use NegPrompt [14] and IDPrompt [15]. These methods involve additional training on OOD prompts. In NegPrompt, the hyperparameters from prior work are applied, and training is done with 16 shots. On the other hand, IDPrompt has a higher learning cost, therefore, following previous studies, training is conducted with just 1 shot. The average value for each of the few-shot methods was calculated over 3 seeds.
3.2 Evaluation Metrics
For evaluation, we use the area under the receiver operating characteristic curve (AUROC). A higher value indicates better performance, with 50% serving as the baseline.
3.3 Findings on ImageNet-X
[…]
The performance rankings changed across the benchmarks. No single winner emerges as the best across all benchmarks.
Conventional Benchmark ImageNet-X ImageNet-FS-X Wilds-FS-X
methods Common-OOD Hard-OOD iWildCam FMoW
Zero-Shot MCM[5] 90.52 63.72 74.32 68.83 76.42 52.80
GL-MCM[10] 92.08 67.43 75.36 70.79 72.46 52.85
NegLabel [11] 94.83 70.44 72.15 68.50 78.74 53.73
EOE [12] 93.11 67.85 77.07 71.76 62.71 55.45
CLIPN-A [13] 93.76 82.51 77.08 69.81 61.60 55.89
Few-Shot CoOp [32] 92.38 67.54 76.63 69.41 71.35 56.44
LoCoOp [6] 93.49 67.77 73.74 68.40 65.50 55.99
NegPrompt [14] 92.73 76.91 80.95 75.14 57.62 56.33
IDPrompt [15] 90.35 60.43 67.42 62.28 60.18 51.37
F1. The effectiveness of methods can vary as the benchmarks. From Table 2, a comparison between the conventional benchmarks and ImageNet-X reveals that the performance rankings changed across the three benchmarks. No single winner emerges as the best across all distribution shifts. Comparing common-OOD and hard-OOD, hard-OOD showed significantly lower overall performance and changes in ranking. While zero-shot methods utilizing OOD prompts outperform others in common-OOD, the best performance in hard-OOD was achieved by CLIPN-A, followed by the few-shot method NegPrompt. Conversely, in ImageNet-X, NegPrompt outperformed CLIPN-A, unlike in hard-OOD. Additionally, LoCoOp outperformed non-OOD prompt methods (MCM, GL-MCM, CoOp) in common-OOD and hard-OOD but showed lower performance compared to these methods in ImageNet-X, indicating notable ranking shifts.
Fig. 3: Difference in OOD detection performance (AUROC) between ImageNet-X and ImageNet-FS-X for CLIP-based methods, calculated as ImageNet-X minus ImageNet-FS-X results.
3.4 Findings on ImageNet-FS-X
F2. The performance drops due to the impact of covariate shifts. When comparing ImageNet-X and ImageNet-FS-X in Table 2, it is observed that in all methods, the performance significantly drops when covariate shift is introduced. These findings indicate that CLIP, even when pre-trained on large datasets, is influenced by covariate shifts, reflecting inherent biases toward specific distributions. The issue of robustness to covariate shifts in CLIP-based classification has been discussed [34], but it must also be considered in OOD detection.
F3. Few-shot learning induces performance bias due to covariate distribution in training data. The performance drop from ImageNet-X to ImageNet-FS-X for each method is shown in Fig. 3. From the graph, it can be seen that the differences are larger for few-shot methods compared to zero-shot methods. This suggests that training with specific data tends to cause a slight performance bias toward the covariate distribution inherent in the training data.
F4. The method performance ranking remains consistent with constant semantic shift. As shown in Table 2, no significant changes were observed in performance rankings when comparing ImageNet-X and ImageNet-FS-X. Calculating the Spearman rank-order correlation coefficient between the benchmarks resulted in a value of 0.90. This suggests that including both semantic and covariate shifts does not substantially alter the overall ranking of methods compared to evaluations based only on semantic shifts. Consequently, while few-shot prompt learning introduces some bias toward the training data’s covariate distribution, the overall ranking remained largely unchanged.
Table 3: Classification accuracy (%) of representative CLIP-based methods on Wilds-FS-X.
iWildCam FMoW
methods Test-ID Test Test-ID Test
Zero-Shot 7.63 14.68 20.14 22.56
Few-Shot CoOp [32] 40.79 41.44 37.97 39.53
LoCoOp [6] 45.43 43.10 38.57 40.25
NegPrompt [14] 36.24 37.45 35.62 37.89
IDPrompt [15] 9.27 12.38 24.94 27.64
3.5 Findings on Wilds-FS-X
F5. Wilds is a highly challenging benchmark with low classification and OOD detection performance. First, the classification accuracy of the CLIP-based methods in the Wilds-FS-X is shown in Table 3. The accuracy was measured on the ID data within the Wilds-FS-X. According to the result, while few-shot learning improves classification accuracy, the iWildCam and FMoW datasets remain challenging and insufficiently addressed by current methods. Furthermore, as shown in Table 2, there is substantial room for improvement in OOD detection performance. Particularly for FMoW, the performance remains close to the baseline, indicating significant difficulty in distinguishing between ID and OOD data. These results indicate that current CLIP-based methods still need improvement in detecting the subtle semantic shifts required for real-world scenarios in the WILDS datasets.
F6. Improved classification accuracy does not necessarily lead to better OOD detection performance. From Table 2 and Table 3, the ranking of OOD detection performance differs from that of classification accuracy. In particular, for iWildCam, classification accuracy improves significantly with few-shot compared to zero-shot. However, in terms of OOD detection, zero-shot methods surpassed few-shot methods in OOD detection, as additional learning improves classification accuracy but hinders detecting subtle semantic shifts between ID and OOD data. These findings indicate that classification and OOD detection on real-world application data remain challenging and require further improvement. Notably, higher classification accuracy does not always mean better OOD detection, emphasizing the need for integrated evaluation of both metrics.
4 Conclusion
In this paper, we proposed three challenging and practical OOD detection benchmarks to support a precise understanding of method characteristics and simulations reflecting real-world conditions. Our experiments on CLIP-based OOD detection methods showed significant performance drops under covariate shifts and real-world-like datasets, revealing room for improvement. Future work should focus on evaluating methods under more challenging, real-world conditions using our proposed benchmarks to identify limitations and guide improvements, enabling broader applicability of OOD detection methods.
ACKNOWLEDGEMENT
This work was partially supported by JST JPMJCR22U4 and JSPS 25H01164.
Appendix
Appendix A Experimental Detail
In this section, we describe the details of each CLIP-based out-of-distribution (OOD) detection method employed in our experiments.
A.1 Zero-shot OOD detection
MCM and GL-MCM. We utilized the pre-trained CLIP-ViT-B/16 model (https://github.com/openai/CLIP). The same pre-trained CLIP model was consistently used across the other methods as well.
NegLabel. For the list of additional labels, we adopted the label set provided by previous studies.
EOE. This method uses a large language model (LLM) to generate additional labels. We employed gpt-3.5-turbo-16k for this purpose. Following prior work, which suggests designing LLM prompts based on visual similarity to produce suitable outlier labels, we used two types of prompts: the Far OOD prompt, which generates domain-level labels for Common-OOD evaluation; and the Near OOD prompt, which generates fine-grained labels for Hard-OOD and our proposed benchmark.
CLIPN. We employed the “no” Text Encoder provided by previous work as a pre-trained model. Among the two proposed OOD scoring algorithms—the Competing-to-win Algorithm and the Agreeing-to-differ Algorithm—we chose the latter (CLIPN-A), as it has been reported to deliver superior detection performance.
A.2 Few-shot OOD detection
The results for each few-shot method were averaged over 3 random seeds.
CoOp and LoCoOp. We used the pre-trained CLIP-ViT-B/16 model and trained with 16 shots, following the same hyperparameters as in prior work (e.g.training epochs = 50, learning rate = 0.002, batch size = 32, token length = 16).
NegPrompt. Prompt learning was conducted in two stages: first, positive prompts were trained for 100 epochs; then, with the positive prompts frozen, negative prompts were trained for 10 epochs. The loss weights ($\beta$ = 0.1, $\gamma$ = 0.05) were taken from previous work. Training was performed with 16 shots. Since the learning rate and batch size were not explicitly stated, we set them to 0.002 and 32, respectively.
IDPrompt. Due to its higher training cost, training was conducted with only 1 shot, as in prior studies. Each image was randomly cropped into 256 patterns, and the top 32 were used for training. Other hyperparameters (e.g.training epochs = 3, learning rate = 0.005, batch size = 1) also followed previous implementations.
Appendix B OOD Score examples
In this section, we present the out-of-distribution (OOD) score distributions for the benchmarks along with several example samples.
Fig. A: OOD score distribution and sample scores in ImageNet-X.
B.1 ImageNet-X
Fig. A presents the OOD score distributions and sample scores in ImageNet-X. The distributions of MCM, CoOp, and NegPrompt are shown as representative examples of CLIP-based methods. Furthermore, one representative sample from both the ID and OOD data is provided as an example.
A comparison of the distributions and sample scores reveals that detection performance varies significantly across methods, even when applied to the same data. For example, with NegPrompt, both samples are correctly identified. However, with MCM and CoOp, distinguishing between the two images as ID and OOD becomes challenging. Notably, CoOp assigns a lower OOD score to (i) than to (ii), resulting in reversed detection outcomes compared to NegPrompt, highlighting substantial differences in performance between the methods.
B.2 ImageNet-FS-X
Fig. B: OOD score distribution and sample scores in ImageNet-FS-X. The samples belong to the “piano” label and are part of the ID data.
Fig. B shows the OOD score distribution and sample score examples in ImageNet-FS-X. The samples are drawn from ImageNet-FS-X, consisting of (i) ImageNet-1k, (ii) ImageNet-V2, (iii) ImageNet-R, and (iv) ImageNet-C, with examples selected based on identical labels across these datasets. These samples belong to the “piano” label and are part of the ID data.
From the sample scores, covariate-shifted ID data is shown to be detectable to some extent, even relative to Training ID. Notably, NegPrompt, which performed best on ImageNet-FS-X, consistently detects all samples as ID with high confidence. Conversely, methods like MCM and CoOp struggle to classify (ii), (iii), and (iv) as ID, though (i) is accurately detected. Similar to ImageNet-X, detection performance varies by method and image. Additionally, while MCM and NegPrompt show scores decreasing in the order of (i), (ii), (iii), and (iv), CoOp assigns a higher score to (iii) than (ii), highlighting differences in recognition order among methods.
Next, we examine the impact of covariate shifts. First, in terms of score distribution, the distributions of ID and OOD in ImageNet-FS-X are closer compared to ImageNet-X. Specifically, the ID data distribution exhibits greater variance, suggesting that covariate shifts cause some ID samples to have scores closer to OOD data. Additionally, in the sample examples, (iv) is an image with noise added to (i). Comparing (i) and (iv), we see that (iv) has a significantly lower score than (i), making detection more difficult. These results demonstrate that differences in style and quality substantially affect OOD detection performance.
Appendix C Covariate shifts analysis
In this section, we compare the OOD detection performance of the source datasets in the full-spectrum benchmarks. This analysis enables further insights into performance variations caused by the covariate shifts and the biases introduced by few-shot learning.
[…]
C.1 ImageNet-FS-X
Fig. C presents the OOD detection performance across source datasets in ImageNet-FS-X. Here, among the variant datasets of ImageNet, ImageNet-V2 represents resampling bias, ImageNet-R reflects style changes, and ImageNet-C serves as an example of image corruption. In all methods, ImageNet-1k achieves the highest performance, followed by ImageNet-V2, ImageNet-R and ImageNet-C. Notably, the AUROC for ImageNet-C is more than 20% lower than ImageNet-1k. These results indicate that CLIP-based methods are sensitive to changes in the covariate distribution of the data, resulting in significant performance variations.
[…]
C.2 Wilds-FS-X
The OOD detection performance for each dataset group, Test and Test-ID, in Wilds-FS-X is shown in Fig. D. First, in zero-shot methods, all datasets demonstrate better performance on the covariate-shifted ID data, Test, compared to the training ID data, Test-ID. This result indicates that the CLIP weights are inherently biased toward achieving higher performance on Test. Furthermore, the observed performance variation due to covariate shifts highlights the need to consider robustness against covariate shifts that can occur in real-world scenarios.
On the other hand, in the case of few-shot methods, Test did not always exhibit higher performance than Test-ID. Notably, in iWildCam, methods such as CoOp, LoCoOp, and IDPrompt achieved higher performance on Test-ID compared to Test. These results further confirm that, even in Wilds-FS-X, few-shot methods exhibit a performance bias toward the covariate distribution of the training ID data.
From the analysis, CLIP’s pre-trained knowledge is biased toward specific covariate distributions, affecting OOD detection performance under covariate shifts. Furthermore, across all benchmarks, a performance bias toward the covariate distribution of the training data was observed when using few-shot learning. Future discussions on the robustness of CLIP-based OOD detection methods to covariate shifts should consider both the biases in pre-trained models and the effects of additional training.